OpenStudy (anonymous):

Solve the equation Quantity x plus 1 all over 3 = 2 for x. Please show your work and solution in the space provided.

OpenStudy (texaschic101):

(x + 1)/3 = 2 Is this your problem ?

OpenStudy (anonymous):

I think I got some of this problem done, x + 1 = 2 ---- 3 3 x + 1 = 3 x 2 ----- 3 x+ 1 = 6 -1 -1 x = 5 - - x x

OpenStudy (texaschic101):

(x + 1)/3 = 2 Multiply the entire equation by 3 3(x + 1)/3 = (3)(2) x + 1 = 6 x = 6 - 1 x = 5 lets check it (x + 1)/3 = 2 (5 + 1)/3 = 2 6/3 = 2 2 = 2 (correct) you did it right...x = 5

OpenStudy (anonymous):

c: yay, thank you.

OpenStudy (texaschic101):

anytime :)